OpenStudy (anonymous):

Find the exact value of the given expression. cos(2 tan^-1 12/5)

OpenStudy (anonymous):

you need two things. "double angle formula" and also \[\cos(\tan^{-1}(\frac{12}{5}))\]

OpenStudy (anonymous):



OpenStudy (anonymous):

from the picture we see that \[\cos(\tan^{-1}(\frac{12}{5})=\frac{5}{13}\] and so now this is like finding \[\cos(2\theta)\] knowing that \[\cos(\theta)=\frac{5}{13}\] \[\cos(2\theta)=2\cos^2(\theta)-1\] so use \[\cos(2\theta)=2\times (\frac{5}{13})^2-1\] to find your answer

OpenStudy (anonymous):

From Mathematica:\[\text{TrigFactor}\left( \cos \left(2 \text{ }\tan ^{-1}\left(\frac{12}{5}\right)\right) \right) = -\frac{119}{169} \]The LHS and the RHS converted to a decimal number respectively:\[\text{Cos}\left[2 \text{ ArcTan}\left[\frac{12}{5}\right]\right]\text{ //}N \to -0.704142 \]\[-\frac{119}{169}\text{ //}N\to -0.704142 \]